OpenStudy (andriod09):

A right triangle has one leg that is 10 inches long, and one leg that is 5 inches long. How long is the triangle's hypotenuse?

OpenStudy (anonymous):

Используйте Пифагора, чтобы найти недостающие стороны. A^ 2 + B ^ 2 = C ^ 2

OpenStudy (calculusxy):

Use the Pythagorean theorem. Do a^2+b^2=c^2

OpenStudy (calculusxy):

And then find the square root of c^2.

OpenStudy (anonymous):

\[h = \sqrt{a^2 + b^2}\]

OpenStudy (calculusxy):

a^2+b^2=c^2 10^2+5^2=100+25=125 125^2=15,625 Squared=125

OpenStudy (anonymous):

\[(hypotenuse)^2 =10^2+5^2 =100+25=125\] \[(hypotenuse) =\sqrt{125} =\sqrt{5^2 \times 5} =5 \sqrt{ 5} \]

OpenStudy (calculusxy):

Did you find your answer?

OpenStudy (andriod09):

Oi vey. I don't understand this. I may be smart, but i can't retain new information, so I don't know how to do this. >.>

OpenStudy (calculusxy):

Did you see my reply?